# FFT-based Dynamic Subspace Selection for Low-Rank Adaptive Optimization of Large Language Models

### v2

@@ 5 Related Work
Our approach aims to improve existing optimizers from prior work in the literature, which we group into two categories: optimizers that speed up convergence at the expense of increased FLOPs and optimizers focused on reducing the running time and memory usage of optimizers by compressing the gradient via low-rank matrix factorization.
@@
The most straightforward approach to computing an orthogonal update is to use the $UV^{\top}$ from the SVD decomposition of the momentum matrix. Since SVD is expensive, it is replaced by an iterative procedure called Newton-Schulz Newton-Schulz, which involves computing the odd powers of the momentum matrix up to $5^{th}$ order and multiplying by some carefully chosen constants. While the Newton-Schulz procedure delivers an accurate approximation of $UV^{\top}$, the odd powers in the polynomial involve full-size matrix multiplications. This is in particular difficult for large scale settings, where the full matrices have to be materialized on a GPU before running Newton-Schulz, which increases communication and memory usage. Dion optimizer (Ahn et al., 2025) aims to reduce the communication overhead by using low-rank, orthogonal updates computed via Power-Iteration, a procedure that requires QR decomposition.
@@
LDAdam (Robert et al., 2025) Block Power Iteration 1 error feedback
Dion (Ahn et al., 2025) Power Iteration 1 save to momentum
Trion (this work) DCT any same as Dion
@@ 6 Conclusion and Limitations
We introduced Trion and DCT-AdamW optimizers that use our DCT-based dynamic column selection to replace two techniques used to perform low-rank decomposition, the inaccurate Power-Iteration in Dion and the expensive SVD in FRUGAL/FIRA and showed that our work improves running time, memory usage and recovers the accuracy of the original methods and thus serves as a cheaper alternative to these expensive and inaccurate methods used to perform low-rank decomposition in adaptive gradient methods for both pretraining and finetuning.

### v3

@@ 5 Related Work
Our approach aims to improve existing optimizers from prior work in the literature, which we group into two categories: optimizers that speed up convergence at the expense of increased FLOPs and optimizers focused on reducing the running time and memory usage by compressing the gradient via low-rank matrix factorization.
@@
The most straightforward approach to computing an orthogonal update is to use the $UV^{\top}$ from the SVD decomposition of the momentum matrix. Since SVD is expensive, it is replaced by an iterative procedure called Newton-Schulz, which involves computing the odd powers of the momentum matrix up to $5^{th}$ order and multiplying by some carefully chosen constants. While the Newton-Schulz procedure delivers an accurate approximation of $UV^{\top}$, the odd powers in the polynomial involve full-size matrix multiplications. This is in particular difficult for large scale settings, where the full matrices have to be materialized on a GPU before running Newton-Schulz, which increases communication and memory usage. Dion optimizer (Ahn et al., 2025) aims to reduce the communication overhead by using low-rank, orthogonal updates computed via Power-Iteration, that requires QR decomposition with running time depending on the rank.
@@
LDAdam (Robert et al., 2025) Block Power-Iteration 1 error feedback
Dion (Ahn et al., 2025) Power-Iteration 1 save to momentum
Trion (this work) DCT 1 same as Dion
@@ 6 Conclusion and Limitations
We introduced the Trion and DCT-AdamW optimizers that use our DCT-based dynamic column selection to replace two techniques used to perform low-rank decomposition, that is, the inaccurate Power-Iteration in Dion and the expensive SVD and QR-decomposition in FRUGAL/FIRA/LDAdam. We showed that our work improves running time and memory usage. Moreover, it recovers the accuracy of the original methods and thus serves as a cheaper alternative to these expensive and inaccurate methods used to perform low-rank decomposition in adaptive gradient methods for both pretraining and finetuning.